Using the following national income accounting data, compute (a) GDP, (b) NDP, and (NI. All figures are in billions. Categoryb. NDP = [ billion C. NIE billion.
Answer #1

ANSWER:

A) GDP = Personal consumption expenditure + net private domestic investment + government purchases + net exports = 219.1 + 52.1 + 59.4 + (17.8 - 16.5) = 331.9

B) NDP = GDP - CONSUMPTION OF FIXED CAPITAL = 331.9 - 11.8 = 320.1

C) NI = NDP - TAXES ON PRODUCTION AND IMPORTS - STATISTICAL DISCREPANCY = 320.1 - 14.4 - 0 = 305.7